Solve this pair of simultaneous equations by substitution method:
y - 3x + 1 = 0
3x + 2y = 16

Answer by LinnW(1048)

y - 3x + 1 = 0
3x + 2y = 16
Add ( 3x - 1) to the top equation
y = 3x - 1
Substitute 3x - 1 for y in the second equation
3x + 2(3x - 1) = 16
3x + 6x - 2 = 16
9x - 2 = 16
add 2 to each side
9x = 18
divide each side by 9
x = 2
Substitute 2 for x in y = 3x - 1
y = 3(2) - 1
y = 6 - 1
y = 5